# Focused-DPO: Enhancing Code Generation Through Focused Preference Optimization on Error-Prone Points

### v1

@@ Abstract
Kechi Zhang,  Ge Li [^1] Corresponding authors.,  Jia Li ♂, Yihong Dong, Jia Li ♀, Zhi Jin [^1] Corresponding authors. Key Lab of High Confidence Software Technology (PKU), Ministry of Education School of Computer Science, Peking University, China {zhangkechi,lige,zhijin}@pku.edu.cn
@@ 1 Introduction
We evaluate Focused-DPO using standard benchmarks such as HumanEval(+) Liu et al. (2024), MBPP(+), and LiveCodeBench Jain et al. (2024), and observe significant improvements over existing methods. Even for models like Qwen2.5-Coder, which already have undergone large-scale alignment training, Focused-DPO still achieves a 42.86% relative improvement on extremely hard competition-level problems in LiveCodeBench. The results show notable increases in the generation quality on error-prone points, highlighting Focused-DPO’s effectiveness in enhancing the accuracy of code generation.
@@ 2 Related Work
Large language models (LLMs) have made significant progress in generating code from natural language descriptions, showing great potential for automating software development tasks. Models(GPT-4, 2023; Li et al., 2023; Hui et al., 2024; Guo et al., 2024; Jiang et al., 2024) have demonstrated strong performance, thanks to extensive training on diverse datasets. To further enhance their capabilities, posting training methods like Supervised Fine-Tuning (SFT) Luo et al. (2023); Wei et al. (2023b) and Direct Preference Optimization Hui et al. (2024); Zhang et al. (2024b); Dou et al. (2024); Gee et al. (2024); Zhang et al. (2024a) are commonly applied. Preference optimization approaches focus on aligning model outputs with desired outcomes by prioritizing more favorable responses over less favorable ones. However, existing DPO approaches fail to address one important issue: they do not directly target the most error-prone points in generated code. Errors in these high-impact parts can lead to significant quality and reliability issues in the final output. We aim to address this issue by focusing the preference optimization learning on these error-prone points in the generated code.
@@ 4 Experiment Setup
To explore this, we evaluate Focused-DPO on several widely-used code generation benchmarks, including HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2024), MBPP (Austin et al., 2021), MBPP+, and LiveCodeBench Jain et al. (2024).
@@ 5.2 Main Results (RQ2)
As shown in Table 5, we start from the base model and progressively incorporate the SFT stage Wei et al. (2023b) and the first DPO stage Zhang et al. (2024b). Finally, applying our Focused-DPO leads to the highest pass rates achieved. These results demonstrate that Focused-DPO effectively boosts the performance of models that have already been extensively post-trained and optimized through previous stages. We further evaluate how Focused-DPO enhances the quality of error-prone points in Appendix C.
@@ 6 Conclusion
We propose Focused-DPO, a framework that improves code generation by focusing on error-prone points. These critical parts significantly impact overall program correctness. Focused-DPO improves Direct Preference Optimization by prioritizing these points, using our Error-Point Identification method to create datasets without costly human annotations. Evaluations show Focused-DPO reduces errors and improves code quality, even in heavily post-trained models. This research highlights the benefits of focusing on fine-grained preference optimization in AI-driven software development.
@@ 7 Acknowledgement
This work is primarily based on our prior research and findings in CodeDPO Zhang et al. (2024b). We are deeply grateful to Jingjing Xu, Jing Su, and Jun Zhang for their valuable discussions regarding the preliminary experiments.
@@
Additionally, our dataset construction pipeline includes specific design choices, such as the use of a page-rank mechanism and the identification of error-prone points based on common prefixes and suffixes. Our preliminary experiments suggest that these settings effectively support the performance of Focused-DPO. Detailed discussions on these designs are provided in Appendix A.
Appendix A Discussion
A.1 Error-Prone Points Identification
A.2 Data Scaling For Focused-DPO
@@ Appendix B Dataset Statistics
Appendix C Improvement in Error-Prone Points
We further evaluate how Focused-DPO enhances the quality in error-prone points. Using our validation dataset (Table 8), we measure the model’s performance on these error-prone parts. The generation probability difference between chosen_mid and reject_mid in error-prone points is illustrated in Figure 4 for the Qwen2.5-Coder-Instruct-7B model.
Appendix D Case Studies for Error-Prone Points

### v2

@@ Abstract
Kechi Zhang1,2,  Ge Li1,2∗,  Jia Li3, Yihong Dong1,2, Jia Li1,2, Zhi Jin1,2∗ 1Key Lab of High Confidence Software Technology (PKU), Ministry of Education 2School of Computer Science, Peking University, China 3College of AI, Tsinghua University {zhangkechi,lige,zhijin}@pku.edu.cn
11footnotetext: Ge Li and Zhi Jin are the corresponding authors.
@@ 1 Introduction
We evaluate Focused-DPO using standard benchmarks such as HumanEval(+) Liu et al. (2024a), MBPP(+), and LiveCodeBench Jain et al. (2024), and observe significant improvements over existing methods. Even for models like Qwen2.5-Coder, which already have undergone large-scale alignment training, Focused-DPO still achieves a 42.86% relative improvement on extremely hard competition-level problems in LiveCodeBench. The results show notable increases in the generation quality on error-prone points, highlighting Focused-DPO’s effectiveness in enhancing the accuracy of code generation.
@@ 2 Related Work
Large language models (LLMs) have made significant progress in generating code from natural language descriptions, showing great potential for automating software development tasks Zhang et al. (2024d, c). Models(GPT-4, 2023; Li et al., 2023; Hui et al., 2024; Guo et al., 2024; Jiang et al., 2024) have demonstrated strong performance, thanks to extensive training on diverse datasets. To further enhance their capabilities, posting training methods like Supervised Fine-Tuning (SFT) Luo et al. (2023); Wei et al. (2023b); Zhang et al. (2023) and Direct Preference Optimization Hui et al. (2024); Zhang et al. (2024b); Dou et al. (2024); Gee et al. (2024); Zhang et al. (2024a) are commonly applied. Preference optimization approaches focus on aligning model outputs with desired outcomes by prioritizing more favorable responses over less favorable ones. However, existing DPO approaches fail to address one important issue: they do not directly target the most error-prone points in generated code. Errors in these high-impact parts can lead to significant quality and reliability issues in the final output. We aim to address this issue by focusing the preference optimization learning on these error-prone points in the generated code.
@@ 4 Experiment Setup
To explore this, we evaluate Focused-DPO on several widely-used code generation benchmarks, including HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2024a), MBPP (Austin et al., 2021), MBPP+, and LiveCodeBench Jain et al. (2024).
@@ 5.2 Main Results (RQ2)
As shown in Table 5, we start from the base model and progressively incorporate the SFT stage Wei et al. (2023b) and the first DPO stage Zhang et al. (2024b). Finally, applying our Focused-DPO leads to the highest pass rates achieved. These results demonstrate that Focused-DPO effectively boosts the performance of models that have already been extensively post-trained and optimized through previous stages. We further evaluate how Focused-DPO enhances the quality of error-prone points in Appendix 6.2.
@@ 5.3 Ablation Study (RQ3)
6 Discussions
6.1 Error-Prone Points Identification
6.2 Improvement in Error-Prone Points
We further evaluate how Focused-DPO enhances the quality in error-prone points. Using our validation dataset (Table 8), we measure the model’s performance on these error-prone parts. The generation probability difference between chosen_mid and reject_mid in error-prone points is illustrated in Figure 4 for Qwen2.5-Coder-Instruct-7B model.
@@ 7 Conclusion
We propose Focused-DPO, a framework that improves code generation by focusing on error-prone points. These critical parts significantly impact overall program correctness. Focused-DPO improves Direct Preference Optimization by prioritizing these points, using our Error-Point Identification method to create datasets without costly human annotations. Evaluations show Focused-DPO reduces errors and improves quality, even in heavily post-trained models. This research highlights benefits of focusing on fine-grained preference optimization in AI-driven software development.
@@ 8 Acknowledgement
This research is supported by the National Natural Science Foundation of China under Grant Nos. 62192731, 62192730, 62192733, 62072007, the National Key R&D Program under Grant No. 2023YFB4503801, and the Major Program (JD) of Hubei Province (No.2023BAA024).  [^2] This work is based on our prior findings in CodeDPO Zhang et al. (2024b). We are deeply grateful to Jingjing Xu, Jing Su, and Jun Zhang for their valuable discussions.
@@
Additionally, our dataset construction pipeline includes specific design choices, such as the use of a page-rank mechanism and the identification of error-prone points based on common prefixes and suffixes. Our preliminary experiments suggest that these settings effectively support the performance of Focused-DPO. Detailed discussions on these designs are provided in Section 6.
Appendix A Data Scaling For Focused-DPO
@@ Appendix B Dataset Statistics
Appendix C Scalability and Dataset Independence
We aim for our proposed method to be scalable and minimally dependent on external prompt resources. Inspired by the self-OSS methodology, we envision that both question generation and solution alignment can be achieved through self-alignment and self-improvement using the policy model itself—a key step toward our ultimate goal.
On the other hand, our method is intentionally designed to avoid being constrained by specific datasets, unlike approaches such as Code-Optimise. Our aspiration is to develop a stable and robust pathway for models to continually enhance their capabilities autonomously.
In our paper, Focused-DPO achieves improvements using preference feedback directly from the model itself. This avoids potential risks associated with externally distilled data, such as unintentionally transferring knowledge from a more capable model to a less capable one.
To further validate the effectiveness of Focused-DPO, we plan to extend our evaluation to existing code datasets. We have adopted the sample datasets from CodeContests. We use the provided test cases to verify the solutions and filter out those samples that have no correct solutions. The final sample number in this dataset setting is nearly 4500. The experiment results are shown in the following Table 10.
Model HumanEval HumanEval+ MBPP MBPP+ LiveCodeBench
Qwen2.5-coder-instruct-7B 0.915 0.841 0.828 0.714 0.312
+ CodeContests dataset
Focused-DPO 0.920 0.860 0.844 0.749 0.334
DPO 0.914 0.847 0.833 0.725 0.308
Table 10: Focused-DPO and DPO Results on CodeContests-train
These results demonstrate that Focused-DPO is both dataset-agnostic and scalable. We plan to expand testing across additional datasets like APPs in future work to further validate reproducibility and fair comparisons.
Appendix D Evaluation of $k$’s Impact
We conducted experiments to measure the consistency of error-prone regions identified across sampled code sequences for varying $k$ values, as shown in Figure 11. The Phi coefficient results reveal that moderate $k$ values (e.g., $k=10$) provide the most consistent identification of error-prone areas.
$k$ Value Strategy Correct Incorrect Phi Coefficient
$k=5$ Common Prefix + Chosen Mid 0.5926 0.0935 0.5237
Common Prefix + Reject Mid 0.0132 0.4629 -0.5314
$k=10$ Common Prefix + Chosen Mid 0.6367 0.0911 0.5651
Common Prefix + Reject Mid 0.0116 0.5575 -0.6085
$k=20$ Common Prefix + Chosen Mid 0.6521 0.0975 0.5710
Common Prefix + Reject Mid 0.0102 0.5473 -0.6025
Table 11: Phi Coefficient for different $k$ settings
We further validated the impact of $k$ by conducting training experiments with different values. The results in Figure 12 indicate that moderate $k$ values yield stable performance improvements across benchmarks.
$k$ Value HumanEval HumanEval+ MBPP MBPP+ LiveCodeBench Avg.
$k=5$ 0.920 0.871 0.835 0.748 0.336
$k=10$ 0.927 0.878 0.847 0.762 0.339
$k=20$ 0.927 0.878 0.838 0.751 0.339
Table 12: Benchmark results for different $k$ settings based on Qwen2.5-coder-instruct-7B
Appendix E Suffix in Positive Loss
In the original DPO loss, penalizing inappropriate negative samples tends to significantly harm performance. However, retaining a portion of the positive sample learning component improves the robustness and stability of model training. Several existing works, including Liu et al. (2024b), have analyzed this phenomenon theoretically. We have also empirically verified this observation in our preliminary experiments. The results, presented in Table 13, demonstrate consistent performance improvements.
Model HumanEval HumanEval+ MBPP MBPP+ LiveCodeBench
Qwen2.5-coder-instruct-7B 0.915 0.841 0.828 0.714 0.312
Focused-DPO 0.927 0.878 0.847 0.762 0.339
No Suffix in Positive Reward 0.927 0.872 0.844 0.760 0.334
Table 13: Benchmark Results Comparison for “No Suffix in Positive Reward”
Appendix F Case Studies for Error-Prone Points